I don't understand this at all, some help would be appreciated..
(x − 0.8)(x − 0.5) [4.6]

Answer by chessace(471)

It is a little weird.
My best guess is that it is an explicit un-named function with an argument:
Define f(x)=(x-0.8)(x-0.5)
Evaluate f with/at x = 4.6
f[4.6] = ...(sub that value for the 2 x's and compute).
f(4.6)= 3.8 * 4.1 = 15.58